Belpre Co. issued $50,000 of 7-year, zero-coupon notes on January 1 when the market yield was 4%. The bond agreement stated that compounding was semiannual. The issue price of the notes was $37,894. How much interest will Belpre Co. report on its income statement during the first year?

  1. $1,378
  2. $1,531
  3. $1,516
  4. $1,365
Answer #1
Interest to be reported during the first year.
First six month interest = $37894*4%*6/12 =$757.88
For Second Six Month = ($37894+757.88)*4%*6/12 =$773.038
Total Interest To Be Reported =$757.88+773.038 =$1530.92
=$1531 (Rounded Off )
Correct Option : b